Two rugby players run towards each other and have an inelastic collision. Given the mass 70 kg and 60 kg respectively, and velocity 3 ms-1 and 4 ms-1. Calculate the final velocity and its direction of the respective players. Discuss how a rugby player’s bone develops if continuously performing the weight bearing exercise using Wolff’s Law and bone remodeling theory


1
Expert's answer
2022-02-08T15:10:10-0500

"m_1v_1+m_2v_2=(m_1+m_2)u,"

"u=\frac{m_1v_1+m_2v_2}{m_1+m_2},"

"u=-0.23~\frac ms."